The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: You-Hua Chu, Dominik J. Bomans, and Eva K. Grebel

MCNews@astro.uiuc.edu

http://www.astro.uiuc.edu/mcnews/MCNews.html

Number: 6 August 19, 1996

Table of Contents

News and Views..............1 Paper Abstracts..............2

News and Views

IAU Symposium on the Magellanic Clouds

The suggestion of an IAU Symposium on the Magellanic Clouds presented in the last MCNews was well-received and prompted several serious suggestions. We are now very pleased to report that we have a host for the meeting!

Dr. Jim Hesser (Dominion Astrophysical Observatory) and his colleagues have been seeking for some time an appropriate meeting that could, in part, acknowledge Dr. Sidney van den Bergh’s many achievements. A year ago Sidney decided to officially retire (i.e., go onto emeritus status but still remain scientifically active) in March, 1998. Consequently, the DAO have been thinking that a meeting in 1998 Northern summer in Victoria would provide an ideal setting to honor Sidney. By far the preferred way to show appreciation to Sidney for his many contributions is through a first-class meeting on a topic that is undergoing major advancement, i.e., Magellanic Cloud research.

Therefore, an offer has been made for the DAO and University of Victoria (UVic) to co-host a Magellanic Clouds meeting, provide the Local Organization Committee, and participate in the Scientific Organization Committee. The meeting could be on the UVic campus, which has the Lam Auditorium (seats 325) and the Recital Hall (seats 220) available at zero or minimal cost. UVic student residences offer standard B&B service in summer months at reasonable prices. DAO may also be able to provide some financial support.

We are very grateful to Jim and his colleagues for their efforts which led to this development. Victoria, located on beautiful Vancouver Island with numerous attractions, will provide an excellent site for the meeting.

A successful meeting results from a combination of many factors. A strong Scientific Organizing Committee is required. The SOC should be chaired by someone who is broadly knowledgeable in Magellanic Cloud research, has a vision of the manifest interactions between multi-wavelength studies of the Magellanic Clouds, and an understanding of larger issues relating to the role of the Clouds in the Local Group and beyond.

The next step will be to make a proposal to the IAU for consideration of this meeting as an IAU Symposium. We would welcome volunteers and suggestions from MCNews readers for people to help with this proposal. A dedicated e-mail account for the symposium will be established in the future. For the time being, please contact mcnews@astro.uiuc.edu, if you would like to help or make suggestions. Updates and progress of the symposium will be posted in the MC News web page at http://www.astro.uiuc.edu/mcnews/MCNews.html.

You-Hua Chu, Jay Gallagher, and Dominik Bomans (with the help of Jim Hesser)

Improved World Wide Web pages of MCnews

Recently, Eva K. Grebel has joined our editorial team. This lead to a complete overhaul of the existing MCNews webpages and to the addition of new pages such as the page on meetings of interest for Magellanic Cloud researchers, see http://www.astro.uiuc.edu/mcnews/MCmeetings.html. We have also established a European mirror of MCNews to make access to our pages easier and faster for European users. The mirror site is located at http://www.astro.uni-wuerzburg.de/∼grebel/MCNews.html. Please check out our webpages!

Paper Abstracts

Molecular abundances in the Magellanic Clouds: II. Deuterated species in the LMC

Y.-N. Chin^{1,2}, C. Henkel^{3}, T.J. Millar^{4}, J.B. Whiteoak^{5}, R. Mauersberger^{6}

1 Institute of Astronomy and Astrophysics, Academica Sinica, P.O.Box 1-87 Nankang, 115 Taipei, Taiwan 2 Radioastronomisches Institut der Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 3 Max-Planck-Institut für Radioastronomie, Auf dem Hügel 69, D-53121 Bonn, Germany 4 Department of Physics, UMIST, P O Box 88, Manchester M60 1QD, United Kingdom 5 Paul Wild Observatory, Australia Telescope National Facility, CSIRO, Locked Bag 194, Narrabri NSW 2390, Australia 6 Steward Observatory, The University of Arizona, Tucson, AZ 85721, U.S.A.

The first definite discoveries of extragalactic deuterium are reported. DCO^{+} has been detected in three and DCN has been measured in one star-forming region of the Large Magellanic Cloud (LMC). While the HCO^{+}/DCO^{+} abundance ratios are found to be 19 ± 3, 24 ± 4, and 67 ± 18 for N113, N44BC and N159HW, respectively, a HCN/DCN abundance ratio of 23 ± 5 is obtained for N113. These results are consistent with a gas temperature of about 20 K and a D/H ratio of about 1.5 × 10^{−5}, consistent with that observed in the Galaxy. If the cloud temperature is closer to 30 K, then a D/H ratio is required to be up to an order of magnitude larger. Because this ratio provides a lower limit to the primordial D/H ratio, it indicates that the baryon mass density alone is unable to close the universe. Accepted by Astronomy and Astrophysics Letter For preprints, contact einmann@biaa21.biaa.sinica.edu.tw

Supernova Remnants in the Magellanic Clouds I: The Colliding Remnants DEM L 316

Rosa Murphy Williams^{1}, You-Hua Chu^{1}, John R. Dickel^{1}, Ross Beyer^{1}, Robert Petre^{2},

R. Chris Smith^{3} and Douglas K. Milne^{4}

1 Astronomy Department, University of Illinois at Urbana-Champaign, Urbana, IL 61801 2 NASA/GSFC code 666, Greenbelt, MD 20771 3 Astronomy Department, University of Michigan, Ann Arbor, MI 48109-1090 4 Australia Telescope National Facility, CSIRO, Box 76, Epping, New South Wales 2121

The Large Magellanic Cloud (LMC) supernova remnant (SNR) DEM L 316 is of considerable interest, as it shows two overlapping shells at radio and optical wavelengths. This morphology suggests that DEM L 316 is in fact two SNRs, which may or may not be spatially connected. We have observed the DEM L 316 remnants at optical, radio, and X-ray wavelengths, in order to address the question of whether the SNRs are interacting, and if so, whether a tunnel might have formed between the remnants, allowing them to equilibrate. Besides the connecting shell morphologies at optical and radio wavelengths, our observations show enhanced O III emission at the juncture between the shells; enhanced X-ray emission from one shell at that juncture; and a change in the structure of the magnetic field near the interaction region. These factors strongly imply that the two remnants are indeed colliding. In addition, we have done a detailed study on the physical properties of the remnants, including an examination of thermal and kinetic energies, and of the balance of thermal and magnetic pressures. Submitted to The Astrophysical Journal For preprints, contact rosanina@astro.uiuc.edu

LH 72, the Center of Constellation III

Knut A.G. Olsen^{1}, Paul W. Hodge^{1}, Eric M. Wilcots^{2}, and Lora Pastwick^{3}

1 University of Washington, Dept. of Astronomy, Seattle, WA 98195 2 University of Wisconsin, Dept. of Astronomy, Madison, WI 53706 3 University of Wyoming, Dept. of Physics and Astronomy, Laramie, WY 82701

Ultraviolet spectroscopy with the IUE satellite is combined with optical photometry to study the brightest blue members of the LMC OB association LH 72. LH 72 lies near the geometric center of Shapley’s Constellation III, a complex of gas and stars that makes up one of the largest known superbubbles. Ages for the stars range from 5 million to 15 million years, with ages increasing from the south to the north end of the association. The youngest stars are associated with the brightest HII emission and have the largest reddening. This pattern of apparent sequential star formation is contrary to the expanding model proposed by Dopita, Mathewson, & Ford (1985), in which the entire superbubble, more than an order of magnitude larger in size than LH 72, evolves in about the same amount of time. Our data, however, are consistent with an alternative model (Reid, Mould, & Thompson 1987) in which star formation proceeds locally after the initial trigger from the superbubble shock. Accepted by The Astrophysical Journal For preprints, contact olsen@astro.washington.edu http://www.astro.washington.edu/olsen/research.html

The MACHO Project LMC Variable Star Inventory IV: Multimode RR Lyrae Stars, Distance to the LMC and Age of the Oldest Stars

C. Alcock^{1,2}, R.A. Allsman^{3}, D. Alves^{1,2,4}, T.S. Axelrod^{5}, A.C. Becker^{6},
D.P. Bennett^{1,2,4,7}, K.H. Cook^{1,2}, K.C. Freeman^{5}, K. Griest^{2,8}, J. Guern^{2,8},
M.J. Lehner^{2,8}, S.L. Marshall^{1,2}, D. Minniti^{1,2}, B.A. Peterson^{5}, M.R. Pratt^{2,6,9},
P.J. Quinn^{10}, A.W.Rodgers^{5}, W. Sutherland^{11}, D.L. Welch^{12}

(The MACHO Collaboration)

1 Lawrence Livermore National Laboratory, Livermore, CA 94550 2 Center for Particle Astrophysics, University of California, Berkely, CA 94720 3 Supercomputing Facility, Australia National University, Canberra, ACT 0200, Australia 4 Dept. of Physics, University of California, Davies, CA 95616 5 Mt. Stromlo and Siding Spring Observatories, Australian National University, Weston Creek, ACT 2611, Australia 6 Depts. of Astronomy and Physics, University of Washington, Seattle, WA 98195 7 Physics Dept., University of Notre Dame, Notre Dame, IN 46556 8 Dept. of Physics, University of California, San Diego, La Jolla, CA 92093 9 Dept. of Physics, University of California, Santa Barbara, CA 93106 10 European Southern Observatory, Karl-Schwarzschild Str. 2, D-85748 Garching, Germany 11 Dept. of Physics, University of Oxford, Oxford OX1 3RH, U.K. 12 Dept. of Physics & Astronomy, McMaster University, Hamilton, Ontario, L8S 4M1 Canada

We report the discovery of 73 double-mode RR Lyrae (RRd) stars in fields near the bar of the LMC. The stars are detected among the MACHO database of short-period variables that currently contains about 7900 RR Lyrae stars. Fundamental periods (P_{0}) for these stars are found in the range 0.46-0.55 days and first overtone-to-fundamental period ratios are found to be in the range 0.742 < P_{1}/P_{0} < 0.748. A significant fraction of our current sample have period ratios smaller than any previously discovered RRd variables. We present mean magnitudes, colors, and lightcurve properties for all LMC RRd stars detected to date. The range in period ratios is unexpectedly large. We present a determination of absolute magnitudes for these stars based primarily on pulsation theory and the assumption that all observed stars are at the fundamental blue edge (FBE) of the instability strip. Comparison of the calibrated MACHO V and R_{KC} photometry with these derived absolute magnitudes yields an absorption-corrected distance modulus to the LMC of 18.57 ± 0.19 mag which is in good agreement with that found (18.5) through comparison of galactic and LMC Cepheids. Adopting this luminosity calibration, we derive an increase in the distance modulus, and thus a reduction in the age found via isochrone fitting for M15 of about 33% and discuss the implications for cosmology. Submitted to The Astrophysical Journal For preprints, contact alcock@igpp.llnl.gov

Variables and the Magellanic Clouds

Patricia Whitelock

South African Astronomical Observatory, PO Box 9, 7935 Observatory, South Africa

This paper reviews our knowledge of the variable stars which are found in the Magellanic Clouds, and which are of interest in the context of microlensing surveys. These include eclipsing binaries, R CrB stars, β Cephei stars, and red-giant variables. The discussion concentrates on red variables, since this class contains by far the largest number of variable stars which will be investigated by the microlensing surveys. The group encompasses luminous AGB stars, Miras, semiregular and irregular variables, and low amplitude red variables. The many areas in which the microlensing studies are likely to make major contributions to our understanding of these stars are highlighted. It is suggested that some of the most interesting results will come from combining the microlensing results with those from other surveys. To appear in: Proceedings of the 12^{th} IAP Colloquium on: Variable Stars and the Astrophysical Return of Microlensing Surveys eds. R. Ferlet & J.-P. Maillard For preprints, contact paw@saao.ac.za